OpenStudy (anonymous):

5(x-7)=55 5x-35=55 5x=90 x=18 Which property was used to write the equation in step 2?

OpenStudy (anonymous):

@hughesy123

OpenStudy (anonymous):

Is this multiple choice? If so, can you provide the answers?

OpenStudy (anonymous):

distributive property addition property of equality subtraction property of equality division property of equality

OpenStudy (anonymous):

It is the distributive property :) Thats because they distribute 5 over x and 7.

OpenStudy (anonymous):

-5(a+3)=-55 what is the value of A My guess is 14 but here are the options -14 -8 8 14

OpenStudy (anonymous):

-5(a+3)=-55 -5a+(-15)=-55 -5a=-40 a=8

OpenStudy (anonymous):

1/2(x+6)=18 What is the value of x? 6 7 1/2 14 1/2 30

OpenStudy (anonymous):

my guess is 6 it makes the best since to me

OpenStudy (anonymous):

1/2(x+6)=18 1/2x + 3=18 1/2x=15 x=7.5 or 7 1/2

OpenStudy (anonymous):

2.5x=5.7 Which method should be used to solve for x? distributive property addition property of equality combining like terms division property of equality

OpenStudy (anonymous):

last one plz.!;3

OpenStudy (anonymous):

The correct answer is Division Property, since you would divide 2.5 across the = sign in order to find x.